If the mean and the standard deviation of a continuous random variable that is normally distributed are 19 and 5 respectively, what is the interval that contains 95% of the distribution? *

1
Expert's answer
2022-03-16T04:03:56-0400

Approximately 95% of the data fall within two standard deviations of the mean:

(μ2σ,μ+2σ)=(1925,19+25)=(9,29).(\mu-2\sigma,\mu+2\sigma) =\\(19-2\cdot5,19+2\cdot5)=(9,29).